OpenStudy (anonymous):

Solve: the logarithm of √5 to the base 2, given that the logarithm of 5 to the base 2 = 2.3219

OpenStudy (anonymous):

use the fact that: \[\log_2(5^{\frac{1}{2}}) = \frac{1}{2}\log_2(5)\]

OpenStudy (anonymous):

its 1/2 * 2.3219 = 1.16095